How much money should be invested today in order to have $8340 at the end of 9 years if the interest paid is a nominal rate of 4.25 percent compounded semi-annually?
Answer #1
The requirement is to find the PV of the future
amount of $8,340 receivable after 9 years when
discounted at an half yearly rate of 2.125%.
The number of periods (half years) would be 9*2 = 18.
Thus, the amount to be deposited today (PV) = 8340/1.02125^18 = $   5,712.01